x pounds of candy valued at $3.50 Per pounds is mixed with y pounds of candy valued at $4.30 per pound to produce 10 pounds of a mixture selling for $4 per pound. Fin x and y, the number of pounds of each type
Answer by mananth(16946)

quantity equation
x+y=10
cost equation
3.5x+4.3y = 10*4
multiply by 10
35x+43y=400
..
x+y=10
35x+43y=400
.
multiply the first equation by 35 and subtract the second
35x+35y-35x-43y=350-400
-8y=-50
y=6.25 pounds
So x= 3.75 pounds